Users' Executive Committee
Users' Executive Committee
Fermi Research Alliance, LLC

Meetings and Minutes

Minutes of the 3 April. 2009 UEC Meeting

Users Executive Committee members attending:

Karen Gibson
Dan Hooper
Matthew Jones
Ashutosh Kotwal
Ron Moore
Kevin Pitts
Heather Ray
Lee Sawyer
Mitch Soderberg
Jon Urheim


Roger Dixon - Accelerator Update


Roger Dixon spoke to members of the UEC and gave a report on the accelerator operations.

Dixon: There have been normal operations in the accelerators since the last UEC update on 7 November 2008. A 10-week shutdown is scheduled to begin in June.

The integrated luminosity from the Tevatron has been out-performing highest expectations.

Director Oddone said we would run through FY2010 at the recent all-hands meeting. FY11 running is not determined yet.

The new method of injection into the Tevatron allow the option to inject only a fraction of the available antiproton stash, leaving some antiprotons in the "bank" for the next store.

The main focus for the accelerator is to integrate luminosity. We will continue to optimize and make minor improvements.

For desired accelerator maintenance, we would need two shutdowns. The scheduling of the second shutdown depends on the overall run plan for the Tevatron in FY10 and FY11.

The Accelerator Division is already working on the plan for how Tevatron workforce will transition to other projects after the end of the Tevatron run.

There is no significant impact on the accelerator due to the proposed extended MiniBooNE running. The protons required by MiniBooNE don't get used by anyone else, and MiniBooNE has a spare horn.

(UEC) Could we combine the two proposed shutdowns into one longer shutdown this summer?
Dixon: Probably not. It is difficult to plan maintenance for two years into the future by this summer, with the present uncertainy of the FY11 run plan. Also, CDF and D0 would be opposed to a longer shutdown.

(UEC) Why is NuMI performance below the design curve?
Dixon: This is due to horn problems.


Users' Meeting Update - Lee Sawyer


Banquet caterer is being finalized. Current caterer does not provide Indian food, and this has caused some problems.

Michael Turner will speak at Outreach Meeting on Friday.

Directorate would like to have the public speaker speak about benefits of HEP.


Update from Washington, D.C. - Carole McGuire


Since the last update, the FY 2009 omnibus appropriation bill and economic stimulus package have been completed. There is still $400M of stimulus money outstanding for the DOE Office of Science, including significant money for Fermilab. This will be settled soon. NSF has submitted a plan for its $3B in stimulus funding to OMB and the White House, for which negotations are ongoing. We don't know yet the full amount going to HEP.

FY 2010 budget details may be delayed until early May. We won't know how much the President is requesting for science in his budget by the time of the April UEC visit to Washington, DC. The UEC visit will be timely since it will highlight the need for funding support for HEP as the Administration prepares to make its official request.

Several members of Congress are requesting an 8 percent increase for DOE Office of Science in the FY 2010 budget. Similar activity is ongoing in support of NSF. Carole will send Dear Colleague letters expressing these requests to the UEC.

The House of Representatives and Senate both passed their versions of the FY2010 budget resolutions last night. Both versions assume less than the President's overall level for discretionary appropriated programs due to deficit worries.

(UEC) Does DOE think that its funding is on the America Competes doubling path now?
McGuire: Yes, but we should be careful. We may not see the 8 percent increase in the FY 2010 budget, due to the funding for science in the stimulus package and the FY 2009 omnibus appropriations bill.


Washington, D.C. Trip Update - Dan Hooper


The coordination meeting at the URA office will be at 8:00 AM on Wednesday morning. First congressional meetings should be at 10:00 AM and beyond on Wednesday, and after 9:00 AM on Thursday.

Materials for UEC trip are nearing completion.

Primary appointments should be done in about two weeks. Secondary assignments will begin after that.